OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 2(n - 3) + 5 = 3(n - 1)

OpenStudy (anonymous):

That's what it say's on my homework...

OpenStudy (anonymous):

n=2

OpenStudy (inkyvoyd):

Yup.

OpenStudy (anonymous):

I need the steps...

OpenStudy (inkyvoyd):

Do you really not know how to do this? We can show you :)

OpenStudy (anonymous):

2(n - 3) + 5 = 3(n - 1) 2(2-3)+5=3(2-1) n=2

OpenStudy (anonymous):

that's the only steps, pretty much